# SOOD++: Leveraging Unlabeled Data to Boost Oriented Object Detection

### v1

@@ Abstract
Semi-supervised object detection (SSOD), leveraging unlabeled data to boost object detectors, has become a hot topic recently. However, existing SSOD approaches mainly focus on horizontal objects, leaving multi-oriented objects common in aerial images unexplored. At the same time, the annotation cost of multi-oriented objects is significantly higher than that of their horizontal counterparts. Therefore, in this paper, we propose a simple yet effective Semi-supervised Oriented Object Detection method termed SOOD++. Specifically, we observe that objects from aerial images are usually arbitrary orientations, small scales, and aggregation, which inspires the following core designs: a Simple Instance-aware Dense Sampling (SIDS) strategy is used to generate comprehensive dense pseudo-labels; the Geometry-aware Adaptive Weighting (GAW) loss dynamically modulates the importance of each pair between pseudo-label and corresponding prediction by leveraging the intricate geometric information of aerial objects; we treat aerial images as global layouts and explicitly build the many-to-many relationship between the sets of pseudo-labels and predictions via the proposed Noise-driven Global Consistency (NGC). Extensive experiments conducted on various multi-oriented object datasets under various labeled settings demonstrate the effectiveness of our method. For example, on the DOTA-V1.5 benchmark, the proposed method outperforms previous state-of-the-art (SOTA) by a large margin (+2.92, +2.39, and +2.57 mAP under 10%, 20%, and 30% labeled data settings, respectively) with single-scale training and testing. More importantly, it still improves upon a strong supervised baseline with 70.66 mAP, trained using the full DOTA-V1.5 train-val set, by +1.82 mAP, resulting in a 72.48 mAP, pushing the new state-of-the-art. Moreover, our method demonstrates stable generalization ability across different oriented detectors, even for multi-view oriented 3D object detectors. The code will be made available.
@@
Oriented object detection, Multi-oriented object detection, Aerial scenes, Semi-supervised learning
@@ 1 Introduction
Sufficient labeled data is essential to achieve satisfactory performance for fully-supervised object detection. However, the data labeling process is expensive and laborious. To alleviate this problem, many Semi-Supervised Object Detection (SSOD) methods [39, 64, 76, 28], aiming to learn from labeled data as well as easy-to-obtain unlabeled data, have been proposed recently. By leveraging the potentially useful information from the unlabeled data, the SSOD methods can achieve promising improvement compared with the supervised baselines, i.e., methods only trained from the limited labeled data.
@@
The recent advanced SSOD methods [39, 64, 76, 28] mainly focus on detecting objects with horizontal bounding boxes in general scenes. However, in more complex scenes, notably aerial scenes, we need to utilize oriented bounding boxes to precisely describe the objects. The complexity of annotating objects with oriented bounding boxes significantly surpasses that of their horizontal counterparts, e.g., the annotation process for oriented bounding boxes incurs an approximate 36.5% increase in costs [^1] The cost of annotating an oriented box is approximately 36.5% more expensive than a horizontal box (\$86 vs. \$63 per 1k) https://cloud.google.com/ai-platform/data-labeling/pricing. Thus, considering the higher annotation cost of oriented boxes, semi-supervised oriented object detection is worth studying.
@@
Figure 1: (a) The comparison of general scenes and aerial scenes. Different from the general objects, the objects from aerial scenes are arbitrary orientations, small, and dense (aggregation). Correspondingly, the annotation cost of aerial objects is higher than that of general scenes. (b) The proposed SOOD++ outperforms the supervised baseline and SOTA methods [35, 76] by a large margin on the DOTA-V1.5 benchmark (val set) under various settings.
@@
Compared with the general scenes, the distinguishing characteristics of objects from aerial scenes (referred to as aerial objects) are mainly from three aspects: arbitrary orientations, small, and dense, as shown in Fig. 1(a). The mainstream SSOD methods are based on the pseudo-labeling framework [5, 55, 54, 64] consisting of the teacher and student models. The teacher model operates as an Exponential Moving Average (EMA) of the student model, aggregated over historical training iterations. It aims to generate pseudo-labels for images within the unlabeled dataset, thereby facilitating the student model’s learning from a combination of labeled and unlabeled data.
@@
However, through pilot experiments, we find that extending the SSOD framework to oriented object detection is a highly non-trivial problem. We think the following two aspects need to be considered: 1) First, the geometry information (e.g., orientation and aspect ratio) is an essential cue for representing the multi-oriented objects. How to use the geometry information for the interaction between teacher and student in a semi-supervised framework is crucial. In other words, when applying SSOD methods in such a scenario, it is necessary to consider the geometry information of pseudo-labels and predictions. 2) Second, aerial objects are often dense and regularly distributed in an image. These properties lead to a challenge for obtaining accurate pseudo-labels from the teacher, resulting in inaccurate knowledge transfer towards the student and damaging the learning effectiveness within a teacher-student framework. A simple yet effective solution is to consider the global perspective. The global appearance of pseudo-labels and predictions can naturally reflect their patterns and suppress local noise, benefiting a more consistent and steady learning process.
@@
Based on these observations and analysis, in this paper, we propose a practical Semi-supervised Oriented Object Detection method termed SOOD++. SOOD++ is built upon the dense pseudo-labeling framework, where the pseudo-labels are filtered from the pixel-wise predictions (including box coordinates and confidence scores). The key designs are from three aspects: a Simple Instance-aware Dense Sampling (SIDS) strategy, Geometry-aware Adaptive Weighting (GAW) loss, and Noise-driven Global Consistency (NGC). The SIDS is used to construct comprehensive dense pseudo-labels from the teacher, similar to pre-processing. GAW and NGC aim to measure the discrepancy between the teacher and student from a one-to-one and many-to-many (global) perspective, respectively.
@@
Specifically, existing dense pseudo-labeling methods [76, 35] directly sample the fixed number of pseudo-labels from the predicted maps, which easily overlooks the dense, small, and ambiguous objects that usually exist in aerial scenes. To construct more comprehensive pseudo-labels, we propose a Simple Instance-aware Dense Sampling (SIDS) strategy to sample the dynamic number of easy and hard cases (e.g., objects intermingled with the background) from the teacher’s output. Second, considering the pseudo-label-prediction pairs are not equally informative, we propose the Geometry-aware Adaptive Weighting (GAW) loss. It utilizes the intrinsic geometry information (e.g., orientation gap and aspect ratio) of each teacher-student pair, reflecting the difficulty of the sample by weighting the corresponding loss dynamically. In this way, we can softly pick those more useful supervision signals to guide the learning of the student.
@@
Furthermore, we innovatively treat the outputs of the teacher and student as two independent global distributions. These distributions implicitly reflect the characteristics of the overall scenario, such as object density and spatial layout across the image. This naturally inspires us to construct global constraints, leading to the proposal of Noise-driven Global Consistency (NGC). Specifically, we first add random noise to disturb the teacher and student outputs (global distributions). Then, We use optimal transport (OT) to align the original teacher’s distribution with the original student’s distribution and similarly align the disturbed teacher’s distribution with the disturbed student’s distribution. Besides, the transport plans from the former two alignments reflect the intricate relationships within the distributions, so we propose aligning them to further evaluate noise impact and provide auxiliary guidance for the model. Applying such multi-perspective alignments will increase the models’ tolerance to minor input variations, enabling them to better capture essential features of the data rather than relying excessively on noise or specific details in the unlabeled data. By promoting global consistency, our NGC not only relieves the negative effect of some inaccuracy pseudo-labels in the unsupervised training stage but also encourages a holistic learning approach, considering objects’ spatial and relational dynamics.
@@
Extensive experiments conducted on four oriented object detection datasets in various data settings demonstrate the effectiveness of our methods. For example, as shown in Fig. 1(b), on the challenging large-scale DOTA-V1.5 (val set), we achieve 50.48, 57.44, and 61.51 mAP under 10%, 20%, and 30% labeled data settings, significantly surpassing the state-of-the-art SSOD method [35] (using the same supervised baseline) by 2.92, 2.39, and 2.57 mAP, respectively. More importantly, it can still improve a 70.66 mAP strong supervised baseline trained using the full DOTA-V1.5 train-val set by +1.82 mAP, resulting in 72.48 mAP on the DOTA-V1.5 (test set) with single-scale training and testing. Moreover, we also achieve consistent improvement on the DOTA-V1.0, HRSC2016, and even nuScenes dataset (a representative multi-view oriented 3D object detection dataset).
@@
This paper is an extended version of our conference paper, published in CVPR 2023 [23], where we make the following new contributions: 1) We propose a Simple Instance-aware Dense Sampling strategy, which not only considers the easily obtained dense pseudo-labels from the foreground but also mine the hard case merged with the background. This design significantly improves the quality of dense pseudo-labels. 2) We propose Geometry-aware Adaptive Weighting (GAW) loss, which leverages the intrinsic geometry information (e.g., orientation gap and aspect ratios) of aerial objects to dynamic weight the importance of a given teacher-student pair. 3) We propose Noise-driven Global Consistency (NGC) to achieve multi-perspective global alignments, encompassing the alignments between the original teacher and original student, the disturbed teacher and disturbed student, and the transport plans generated from the former alignments. 4) Through the above substantial technical improvements, our SOOD++ significantly outperforms its conference version SOOD [23] by +1.85, +1.86, and +2.28 mAP on the DOTA-V1.5 dataset under the 10%, 20%, and 30% labeled settings, respectively. 5) We have made significant improvements in the quality of the manuscript in various aspects, e.g., conducted experiments on more benchmarks (even 3D scenes), more profound insights into performance improvements, more thorough ablation studies, added more meaningful illustrative figures, and reorganized the contents and tables to better illustrate the paper, etc.
@@ 2 Related works
Semi-Supervised Object Detection.
@@
In recent years, semi-supervised learning (SSL)[52, 1] has demonstrated remarkable performance in image classification. These methods exploit unlabeled data through various techniques such as pseudo-labeling[27, 18, 61, 29], consistency regularization [55, 60, 1], data augmentation [51, 6], and adversarial training [44]. Unlike semi-supervised image classification, semi-supervised object detection (SSOD) necessitates instance-level predictions and the additional task of bounding box regression, thereby increasing its complexity. In [48, 77, 53, 25], pseudo-labels are generated by combining predictions from various augmentation.
@@
Subsequent studies [39, 54, 64, 67] have incorporated EMA from Mean Teacher [55] to update the teacher model after each training iteration. ISMT [67] integrates current pseudo-labels with historical ones. Unbiased Teacher [39] addresses the class-imbalance issue by replacing the cross-entropy loss with focal loss [34]. Humble Teacher [54] utilizes soft pseudo-labels as the student’s training targets, enabling the student to distill richer information from the teacher. Soft Teacher [64] adaptively weights the loss of each pseudo-box based on classification scores and introduces box jittering to select reliable pseudo-labels. Unbiased Teacher v2 [40] employs an anchor-free detector and uses uncertainty predictions to select pseudo-labels for the regression branch. Dense Teacher [76] replaces post-processed instance-level pseudo-labels with dense pixel-level pseudo-labels, effectively eliminating the influence of post-processing hyper-parameters. To ensure both label-level and feature-level consistency, PseCo [28] proposes Multi-view Scale-invariant Learning, which aligns shifted feature pyramids between two images with identical content but different scales. Consistent Teacher [57] combines adaptive anchor assignment, a 3D feature alignment module, and a Gaussian Mixture Model to reduce matching and feature inconsistencies during training. MixTeacher [36] explicitly enhances the quality of pseudo-labels to handle the scale variation of objects. The Semi-DETR [75] method is specifically designed for DETR-based detectors, mitigating the training inefficiencies caused by bipartite matching with noisy pseudo-labels. To address assignment ambiguity, ARSL [35] proposes combining Joint-Confidence Estimation and Task-Separation Assignment strategies.
@@
However, all the methods mentioned above are designed for the general scenes. How to unleash the potential of the semi-supervised framework in aerial scenes has yet to be fully explored. This paper aims to fill this blank and offer a solid baseline for future research.
Oriented Object Detection.
@@
Unlike general object detectors [17, 50, 37, 49], representing objects with Horizontal Bounding Boxes (HBBs), oriented object detectors use Oriented Bounding Boxes (OBBs) to capture the orientation of objects, which is practical for detecting aerial objects. In recent years, numerous methods [70, 32, 43, 10, 65, 42, 14, 73, 7, 8] have been developed to enhance the performance of oriented object detection. For instance, CSL [69] addresses the out-of-bound issue by transforming the angle regression problem into a classification task. R3Det [70] improves detection speed by predicting HBBs in the first stage and then aligning features in the second stage to detect oriented objects. Oriented R-CNN [62] introduces a simplified multi-oriented region proposal network and uses midpoint offsets to represent arbitrarily oriented objects. ReDet [21] incorporates rotation-equivariant networks into the detector to extract rotation-equivariant features. Oriented RepPoints [31] incorporates a quality assessment module and a sample assignment scheme for adaptive point learning, which helps obtain non-axis features from neighboring objects while ignoring background noise. LSKNet [32] extends the large and selective kernel mechanisms to improve the performance. The recent method COBB [59] employs nine parameters derived from continuous functions based on the outer HBB and OBB area.
@@
The above methods typically focus on the fully supervised paradigm, requiring expensive labeling costs. Thus, several weakly supervised object detectors have been proposed recently, such as point-based methods [42] or HBB-based methods [74, 71]. However, they are either inferior to their fully supervised counterparts or fail to leverage unlabeled data for performance improvement. In this paper, we explore semi-supervised oriented object detection, which reduces annotation costs and boosts detectors, even those well-trained on large-scale labeled data, by utilizing additional unlabeled data.

### v2

@@ Abstract
Semi-supervised object detection (SSOD), leveraging unlabeled data to boost object detectors, has become a hot topic recently. However, existing SSOD approaches mainly focus on horizontal objects, leaving oriented objects common in aerial images unexplored. At the same time, the annotation cost of oriented objects is significantly higher than that of their horizontal counterparts (an approximate 36.5% increase in costs). Therefore, in this paper, we propose a simple yet effective Semi-supervised Oriented Object Detection method termed SOOD++. Specifically, we observe that objects from aerial images usually have arbitrary orientations, small scales, and dense distribution, which inspires the following core designs: a Simple Instance-aware Dense Sampling (SIDS) strategy is used to generate comprehensive dense pseudo-labels; the Geometry-aware Adaptive Weighting (GAW) loss dynamically modulates the importance of each pair between pseudo-label and corresponding prediction by leveraging the intricate geometric information of aerial objects; we treat aerial images as global layouts and explicitly build the many-to-many relationship between the sets of pseudo-labels and predictions via the proposed Noise-driven Global Consistency (NGC). Extensive experiments conducted on various oriented object datasets under various labeled settings demonstrate the effectiveness of our method. For example, on the DOTA-V2.0/DOTA-V1.5 benchmark, the proposed method outperforms previous state-of-the-art (SOTA) by a large margin (+2.90/2.14, +2.16/2.18, and +2.66/2.32) mAP under 10%, 20%, and 30% labeled data settings, respectively, with single-scale training and testing. More importantly, it still improves upon a strong supervised baseline with 70.66 mAP, trained using the full DOTA-V1.5 train-val set, by +1.82 mAP, resulting in a 72.48 mAP, pushing the new state-of-the-art. Moreover, our method demonstrates stable generalization ability across different oriented detectors, even for multi-view oriented 3D object detectors. The project page is at https://dk-liang.github.io/SOODv2/.
@@
Oriented object detection, Aerial scenes, Semi-supervised learning
@@ 1 Introduction
Sufficient labeled data is essential to achieve satisfactory performance for fully-supervised object detection. However, the data labeling process is expensive and laborious. To alleviate this problem, many Semi-Supervised Object Detection (SSOD) methods [1, 2, 3, 4], aiming to learn from labeled data as well as easy-to-obtain unlabeled data, have been proposed recently. By leveraging the potentially useful information from the unlabeled data, the SSOD methods can achieve promising improvements compared with the supervised baselines, i.e., methods only trained from the limited labeled data.
@@
The recent advanced SSOD methods [1, 2, 3, 4] mainly focus on detecting objects with horizontal bounding boxes [^1] Horizontal objects include both horizontally and vertically aligned ones, as they are represented by axis-aligned bounding boxes without rotation. in general scenes. However, in more complex scenes, notably aerial scenes, we need to utilize oriented bounding boxes to precisely describe the objects. The complexity of annotating objects with oriented bounding boxes significantly surpasses that of their horizontal counterparts, e.g., the annotation process for oriented bounding boxes incurs an approximate 36.5% increase in costs [^2] The cost of annotating an oriented box is approximately 36.5% more expensive than a horizontal box (\$86 vs. \$63 per 1k) https://cloud.google.com/ai-platform/data-labeling/pricing. Thus, considering the higher annotation cost of oriented boxes, semi-supervised oriented object detection is worth studying.
@@
Figure 1: (a) The comparison of general scenes and aerial scenes. Different from the general objects, the objects from aerial scenes are arbitrary orientations, small, and dense distribution. Correspondingly, the annotation cost of aerial objects is higher than that of general scenes. (b) The proposed SOOD++ outperforms the supervised baseline and SOTA methods [5, 6, 7] by a large margin on the DOTA-V1.5 benchmark (val set) under various settings.
@@
Compared with general scenes, the distinguishing characteristics of objects from aerial scenes (referred to as aerial objects) are mainly from three aspects: arbitrary orientations, small, and dense distribution, as shown in Fig. 1(a). The mainstream SSOD methods are based on the pseudo-labeling framework [8, 9, 10, 2] consisting of the teacher and student models. The teacher model operates as an Exponential Moving Average (EMA) of the student model, aggregated over historical training iterations. It aims to generate pseudo-labels for images within the unlabeled dataset, thereby facilitating the student model’s learning from a combination of labeled and unlabeled data.
@@
However, through pilot experiments, we find that extending the SSOD framework to oriented object detection is a highly non-trivial problem. We think the following two aspects need to be considered: 1) Geometry information is crucial for representing oriented objects, where the bounding box orientation defines the object’s direction and the aspect ratio represents the width-to-height proportion. In a semi-supervised framework, it is essential to leverage the intrinsic geometric information for effective teacher-student interaction. 2) Aerial objects are often dense and regularly distributed in an image. These properties lead to a challenge for obtaining accurate pseudo-labels from the teacher, resulting in inaccurate knowledge transfer towards the student and damaging the learning effectiveness within a teacher-student framework. A simple yet effective solution is to consider the global perspective. The global appearance of pseudo-labels and predictions can naturally reflect their patterns and suppress local noise, benefiting a more consistent and steady learning process.
@@
Based on these observations and analysis, in this paper, we propose a practical Semi-supervised Oriented Object Detection method termed SOOD++. SOOD++ is built upon the dense pseudo-labeling framework, where the pseudo-labels are filtered from the pixel-wise predictions (including box coordinates and confidence scores). The key designs are from three aspects: a Simple Instance-aware Dense Sampling (SIDS) strategy, Geometry-aware Adaptive Weighting (GAW) loss, and Noise-driven Global Consistency (NGC). The SIDS is used to construct comprehensive dense pseudo-labels from the teacher, similar to pre-processing. GAW and NGC aim to measure and minimize the discrepancy between the teacher and student from an individual-level and global-level perspective, respectively.
@@
Specifically, existing dense pseudo-labeling methods [3, 5] directly sample the fixed number of pseudo-labels from the predicted maps, which easily overlooks the dense, small, and ambiguous objects that usually exist in aerial scenes. To construct more comprehensive pseudo-labels, we propose a Simple Instance-aware Dense Sampling (SIDS) strategy to sample the dynamic number of easy and hard cases (e.g., objects intermingled with the background) from the teacher’s output.
Considering the pseudo-label-prediction pairs are not equally informative, we propose the Geometry-aware Adaptive Weighting (GAW) loss. It utilizes the intrinsic geometry information (e.g., orientation gap and aspect ratio) of each teacher-student pair, reflecting the difficulty of the sample by weighting the corresponding loss dynamically. In this way, we can softly pick those more useful supervision signals to guide the learning of the student.
@@
Furthermore, we innovatively treat the outputs of the teacher and student as two independent global distributions. These distributions implicitly reflect the characteristics of the overall scenario, such as object density and spatial layout across the image. This naturally inspires us to construct global constraints, leading to the proposal of Noise-driven Global Consistency (NGC). Specifically, we first add random noise to disturb the teacher and student outputs (global distributions). Then, we use optimal transport (OT) to align the original teacher’s distribution with the original student’s distribution and similarly align the disturbed teacher’s distribution with the disturbed student’s distribution. Besides, the transport plans from the former two alignments reflect the intricate relationships within the distributions, so we propose aligning them to further evaluate noise impact and provide auxiliary guidance for the model. Applying such multi-perspective alignments will increase the models’ tolerance to minor input variations, enabling them to better capture essential features of the data rather than relying excessively on noise or specific details in the unlabeled data. By promoting global consistency, our NGC not only relieves the negative effect of some inaccurate pseudo-labels in the unsupervised training stage but also encourages a holistic learning approach, considering objects’ spatial relations.
@@
Extensive experiments conducted on six real-world oriented object detection datasets in various data settings demonstrate the effectiveness of our methods. For example, as shown in Fig. 1(b), on the challenging large-scale DOTA-V1.5 dataset (validation set), we achieve 50.48, 57.44, and 61.51 mAP under 10%, 20%, and 30% labeled data settings, significantly surpassing the state-of-the-art SSOD method [6] by 2.14, 2.18, 2.32 mAP, respectively. More importantly, it can still improve a 70.66 mAP strong supervised baseline trained using the full DOTA-V1.5 train-val set by +1.82 mAP, resulting in 72.48 mAP on the DOTA-V1.5 dataset (test set) with single-scale training and testing. We also achieve consistent improvement on the challenging DOTA-V2.0, DOTA-V1.0, HRSC2016, DIOR-R, and even the nuScenes dataset (a representative multi-view oriented 3D object detection dataset).
@@ 2 Related works
2.1 Semi-Supervised Object Detection
@@
In recent years, semi-supervised learning (SSL)[11, 12] has demonstrated remarkable performance in image classification. These methods exploit unlabeled data through various techniques such as pseudo-labeling[13, 14, 15, 16], consistency regularization [9, 17, 12], data augmentation [18, 19], and adversarial training [20]. Unlike semi-supervised image classification, semi-supervised object detection (SSOD) necessitates instance-level predictions and the additional task of bounding box regression, thereby increasing its complexity. In [21, 22, 23, 24], pseudo-labels are generated from various augmentations.
@@
Subsequent studies [1, 10, 2, 25, 26, 4] incorporate Exponential Moving Average (EMA) from Mean Teacher [9] to update the teacher model. ISMT [25] integrates current pseudo-labels with historical ones. Unbiased Teacher [1] addresses the class-imbalance issue by replacing the cross-entropy loss with focal loss [27], and Unbiased Teacher v2 [28] uses uncertainty predictions to select pseudo-labels for the regression branch. Humble Teacher [10] utilizes soft pseudo-labels as the student’s training targets, enabling the student to distill richer information from the teacher. Soft Teacher [2] adaptively weights the loss of each pseudo-box based on classification scores and introduces box jittering to select reliable pseudo-labels. Dense Teacher [3] replaces post-processed instance-level pseudo-labels with dense pixel-level pseudo-labels, effectively eliminating the influence of post-processing hyperparameters. Consistent Teacher [29] combines adaptive anchor assignment and a Gaussian Mixture Model to reduce matching and feature inconsistencies during training. The Semi-DETR [30] method is specifically designed for DETR-based detectors, mitigating the training inefficiencies caused by bipartite matching with noisy pseudo-labels. Based on Semi-DETR, Sparse Semi-DETR [6] improves object query quality with a query refinement module, improving the detection of small and obscured objects. To address assignment ambiguity, ARSL [5] proposes combining Joint-Confidence Estimation and Task-Separation Assignment strategies. TMR-RD-v2 [7] refines model weights dynamically and maintains teacher-student model distinctiveness.
@@
However, all the methods mentioned above are designed for the general scenes. How to unleash the potential of the semi-supervised framework in aerial scenes has yet to be fully explored.
2.2 Oriented Object Detection
@@
Unlike general object detectors [31, 32, 33, 34], representing objects with Horizontal Bounding Boxes (HBBs), oriented object detectors use Oriented Bounding Boxes (OBBs) to capture the orientation of objects, which is practical for detecting aerial objects. In recent years, numerous methods [35, 36, 37, 38, 39, 40, 41, 42, 43, 44] have been developed to enhance the performance of oriented object detection. For instance, CSL [45] addresses the out-of-bound issue by transforming the angle regression problem into a classification task. R3Det [35] improves detection speed by predicting HBBs in the first stage and then aligning features in the second stage to detect oriented objects. Oriented R-CNN [46] introduces a simplified oriented region proposal network and uses midpoint offsets to represent arbitrarily oriented objects. ReDet [47] incorporates rotation-equivariant networks into the detector to extract rotation-equivariant features. Oriented RepPoints [48] incorporates a quality assessment module and a sample assignment scheme for adaptive point learning, which helps obtain non-axis features from neighboring objects while ignoring background noise. LSKNet [36] extends the large and selective kernel mechanisms to improve the performance. The recent method COBB [49] employs nine parameters derived from continuous functions based on the outer HBB and OBB area.
@@
The above methods typically focus on the fully supervised paradigm, requiring expensive labeling costs. Thus, several weakly supervised object detectors have been proposed recently, such as point-based methods [40], horizontal bounding boxes-based methods [50, 51] or hybrid-based methods [52]. For example, H2RBox [50] learns the rotation via self-supervised learning, whose loss measures the consistency of the predicted angles in two different views. Recently, Relational Matching [52] introduces a weak-semi supervised setting, which comprises some fully labeled and some weakly labeled data with a single point. However, these methods are either inferior to their fully supervised counterparts or fail to leverage completely unlabeled data for performance improvement. In this paper, we explore semi-supervised oriented object detection, which reduces annotation costs and boosts detectors, even those well-trained on large-scale labeled data, by utilizing additional unlabeled data.
2.3 Semi-Supervised Oriented Object Detection
Our conference work [53] SOOD is the first to explore the semi-supervised oriented object detection task. It contains two key components: the Rotation-aware Adaptive Weighting (RAW) loss is used to dynamically weigh the pseudo-label-prediction pairs, and the Global Consistency (GC) is designed to establish a many-to-many relationship between the sets of pseudo-labels and predictions.
In this paper, we make the following extensions: 1) We propose a Simple Instance-aware Dense Sampling (SIDS) strategy, which allows us to mine challenging objects that blend into the background, ensuring that both easy and hard cases are adequately represented, significantly improving the quality of pseudo-labels. 2) We refine the RAW loss by further considering the oriented objects’ aspect ratio to adaptively weigh the importance of pseudo-labels-prediction pairs, providing a clearer direction for optimization during the semi-supervised learning. 3) We refine the GC from the single-perspective global alignment to multi-perspective global alignment, which encompasses the alignments between the original teacher and original student, the disturbed teacher and disturbed student, and the transport plans generated from the former alignments. 4) Besides evaluating the widely-used DOTA-V1.5 dataset in the conference version, we conduct experiments on more benchmarks, such as the DOTA-V1.0 and HRSC2016 datasets. Additionally, we perform experiments on semi-supervised oriented 3D object detection on the large-scale nuScenes dataset to demonstrate the generalization of our method.
Through the above substantial technical improvements, our SOOD++ significantly outperforms its conference version SOOD [53] by +2.15/1.85, +1.92/1.86, and +2.33/2.28 mAP on the DOTA-V2.0/DOTA-V1.5 dataset under the 10%, 20%, and 30% labeled settings, respectively. For more details, please refer to Sec. 4 and Sec. 5.3.
Figure 2: An intuitive example illustrating different pseudo-labels: (a) Unlabeled data. (b) Sparse pseudo-label paradigm [2, 28]. (c) Dense sampling based on Top-K [3, 5]. (d) Dense sampling based on foreground (easy case) [53]. (e) The proposed dense sampling strategy incorporates both foreground and background elements.